# Electric Dipole Moments in 5+3 Flavor Weak Effective Theory

### v1

@@ 1 Introduction
Because of the strength of existing and upcoming constraints, it is important to address the question of what existing EDM experiments teach us about CP violation in generic BSM scenarios. In addition, assuming multiple observations of non-zero EDMs in the next generation of experiments, we will be faced with the “inverse problem” of identifying the fundamental mechanism of CP violation from a set of low-energy measurements. Effective Field Theories (EFT) provide the theoretical tools to answer these questions with minimal reliance on specific models of BSM physics Buras:2020xsm . First of all, EFTs such as the Weak EFT (WET) Jenkins:2017jig ; Jenkins:2017dyc allow one to organize all possible flavor-diagonal ($\Delta F=0$) CP-violating operators of a given dimension. Secondly, in WET it is possible to systematically consider loop effects, such as renormalization group evolution and matching corrections induced by heavy quark thresholds, and to connect low-scale observables with physics at the electroweak scale. Here one can match WET onto other EFTs, such as the Standard Model EFT (SMEFT) Buchmuller:1985jz ; Grzadkowski:2010es or the Higgs EFT (HEFT) Buchalla:2013rka ; Buchalla:2017jlu , and study collider phenomenology. This tower of EFTs provides a general and model-independent link between EDMs, other precision flavor observables, and collider experiments, which will be necessary to solve the aforementioned inverse problem.
@@
The advantages of parameterizing EDMs in terms of a minimal set of low-energy operators have long been recognized Pospelov:2005pr ; Khriplovich:1997ga , and there is a vast literature on EDM calculations in specific BSM models (see Pospelov:2005pr ; Engel:2013lsa ; Chupp:2017rkp and references therein). In EFTs, EDMs in the SMEFT at tree level and the matching onto hadronic EFTs were considered in Ref. deVries:2012ab . Beyond the tree level, the contributions to EDMs induced by classes of SMEFT operators, including in the Yukawa Brod:2013cka ; Chien:2015xha ; Buras:2020xsm ; Brod:2022bww ; Bahl:2022yrs ; Fuchs:2020uoc , top Cirigliano:2016nyn ; Fuyuto:2017xup , heavy flavor Ema:2022pmo ; Haisch:2021hcg , Higgs-gauge Dekens:2013zca ; Cirigliano:2019vfc ; Haisch:2019xyi , right-handed charged current Alioli:2017ces and flavor Endo:2019mxw ; Fajfer:2023gie sectors, have been studied in a mostly piecemeal approach. A more systematic analysis of the one-loop contributions to EDMs in SMEFT was carried out in Ref. Kley:2021yhn , but mostly with a focus on the electron and nucleon EDMs. This paper is a first step towards a complete study of $\Delta F=0$ CP violation in EFT extensions of the SM, and its correlations with other sectors of the theory (such as the Higgs and flavor sectors). We focus here on CP-violating (CPV) $\Delta F=0$ WET operators, with $n_{f}=5$ quark flavors and $n_{\ell}=3$ lepton flavors. The main new results of this work are:
@@
  • We derive master formulae for the neutron, proton, and diamagnetic EDMs, and for the precession frequencies in molecular EDM experiments, in terms of the full set of $\Delta F=0$ WET operators at the electroweak scale. The master formulae for $\omega_{\rm HfF}$, and $\omega_{\rm ThO}$, $\omega_{\rm YbF}$ are given in Tables 7, and 15, respectively. Contributions to $d_{\rm Hg}$ are listed in Tables 8 and 9, while the neutron and proton EDMs, $d_{n}$ and $d_{p}$ are given in Tables 10 and 16, respectively.
@@
  • We consider the contributions of operators with heavy quark ($c$ and $b$) and lepton ($\mu$ and $\tau$) flavors, where renormalization group mixing and matching at the heavy fermion thresholds are crucial. In the case of scalar four-quark operators, all operators with two heavy and two light quarks contribute to EDMs at leading logarithmic accuracy. Scalar operators with four heavy quarks do not contribute at the leading log. In this case, we identify next-to-leading logarithmic contributions, arising from one-loop running and one-loop matching at the heavy flavor threshold. Since the renormalization group mixing is mostly driven by QCD, $d_{n}$ and $d_{\rm Hg}$ provide strong constraints even on scalar operators with only heavy quarks.
  • Four-quark vector operators with two heavy and two light quarks also contribute at the next-to-leading log. In this case, the dominant contributions arise from two-loop running and one-loop matching. Four-quark vector operators with two $b$ and two $c$ quarks contribute at the next-to-next-to-leading log. EDM contributions from scalar and vector operators with the same flavors are comparable.
@@
The paper is organized as follows. In Section 2 we introduce the WET basis of $\Delta F=0$ CP-violating operators. In Section 3 we provide expression for the neutron, atomic, and molecular EDMs in terms of WET operators at low energy. In Section 4 we discuss the renormalization group evolution and threshold effects and in Section 5 we provide master formulae for EDMs in terms of WET coefficients at the electroweak scale. In Section 6 we discuss one example of application of the master formulae, EDM constraints on non-standard Higgs couplings. We conclude in Section 7.
@@ 2 Flavor-conserving CP-violation in WET
There are numerous sources of CP-violation in WET at the dim-4, 5, and 6 levels. Since our main goal is to study EDMs, we will restrict ourselves to flavor-conserving WET operators. We can identify two kinds of contributions. Operators constructed out of light fermion flavors, like $u,d$, or $s$ quarks or electrons, generate EDM at the tree level. As we will see, they will receive the strongest constraints. Operators that contain heavier quark and lepton generations can also contribute via loop effects originating either from the renormalization group running or from threshold effects. Below we consider the full list of WET operators with five quark and three lepton flavors that can contribute to EDMs. Very schematically, WET Wilson coefficients at low energy are obtained by solving renormalization group equations, which give rise to expressions of the form
@@
where $g$ here stands for any coupling in WET, and $\mu_{ew}$ is a scale close to the electroweak. The first and second terms in the bracket on the r.h.s of Eq. (1) denote the leading logarithmic (LL) and next-to-leading logarithmic (NLL) series. As we will discuss, the great majority of $\Delta F=0$ CP-violating operator feeds into light flavor operators at LL. These contributions are captured by the solution of the WET one-loop RGEs, which are known Buchalla:1995vs ; Aebischer:2015fzz ; Jenkins:2017dyc . Certain classes of operators, including, for example, vector left-right operators, do not contribute at LL. In these cases, we identify next-to-leading logarithmic (NLL) corrections, arising from 2-loop running and one-loop matching at the heavy quark threshold. A summary of the order in which each WET operator contributes is given in Table 1, where operators contributing at LL, NLL, and NNLL are indicated by no asterisk, or single and double asterisks, respectively. We stress that, while the logarithmic series expansion in Eq. (1) provides useful guidance to organize WET contributions to EDMs, and we will therefore use it throughout the paper, the RGEs involve small couplings (e.g. the QED coupling), for which the suppression from additional loops is not fully offset by the large logarithms. In these cases, it is possible that higher loop corrections to matching and running will be more important than those included here. These can only be captured by pushing the calculations of WET anomalous dimensions and matching corrections to higher order. For the sake of generality, we also include the operators that cannot be generated in SMEFT at leading order, i.e. the ones that violate the hypercharge quantum number.
In the JMS basis complete 1-loop anomalous dimension matrices (AMDs) are known Jenkins:2017jig ; Jenkins:2017dyc Moreover, the complete 1-loop matching conditions of WET onto SMEFT has been recently worked out Jenkins:2017jig in the JMS basis. Given these developments, it is useful to express the EDMs in terms of the Wilson coefficients (WCs) of the JMS basis so that their connection to UV physics can be easily established. The translation of the JMS WCs into several other popular WET bases or vice versa can be numerically carried out using wilson Aebischer:2018bkb program, which we also use to perform LL component of the RG running. Moreover, throughout this paper, we will follow the WCxf conventions for WET and SMEFT Aebischer:2017ugx .
@@ 2.1 Operator Basis for 3+1 Flavors
where $g_{s}$ is the strong coupling constant, $\widetilde{G}^{a}_{\mu\nu}=\frac{1}{2}\varepsilon_{\mu\nu\alpha\beta}G^{a\,\alpha\beta}$. Throughout this paper, we assume that the contribution of the $\bar{\theta}$ term to EDMs is canceled by the Peccei-Quinn (PQ) mechanism Peccei:1977ur ; Peccei:1977hh .
@@ 2.1.1 Dipole Operators
Here $ii$ indicates the flavor index. Down-type type operators having $ii=11,22$, and up-type quarks with $ii=11$ directly contribute to the neutron and atomic EDMs. Dipole operators with $b$ and $c$ quarks contribute via loops. In the leptonic sector, $ii=11$, $22$, and $33$ contribute directly to electron, muon, and $\tau$ EDMs, respectively. Since the limits on $d_{\mu}$ and $d_{\tau}$ are much weaker than on the electron EDM, it is important also to consider the matching and running of heavier leptons into operators leptonic and semileptonic operators that contribute to molecular EDMs.
@@ 2.1.2 Weinberg Operator
The CP-violating Weinberg operator has the form
@@
Here $G_{\mu\nu}$ is the gluon field strength and $f^{abc}$ are the structure constants of $SU(3)_{c}$ group. Note that the WC $\big{[}{C}_{\widetilde{G}}\big{]}$ is real by definition.
@@ 2.1.3 Four-fermion operators with 3+1 Flavors
In this section, we give a complete list of four-fermion WET operators with $n_{f}=3+1$ quark and lepton flavors. All of these operators enter directly into the EDM expressions, hence these are the leading operators. A four-fermion basis in terms of scalar-pseudoscalar and tensor operators was presented in Ref. Buhler:2023gsg , where the one-loop matching to the gradient flow scheme used in Lattice QCD calculations is also provided. Here, to facilitate the renormalization group evolution and the matching to EFTs at the electroweak scale, we adopt the JMS basis and provide a conversion in the Appendix. In the first category, we consider vector operators having $(LL)(RR)$ chiralities of the fermion currents. These are given by
@@
In total, there are 24 four-quark operators with $n_{f}=3$, in agreement with Ref. Buhler:2023gsg . Semi-leptonic operators contribute molecular and atomic EDMs at the tree level
@@
The relevant flavor structures for semi-leptonic operators are
$[O_{ed}^{T,RR}]_{1122}.$
$[{O}_{ee}^{S,RR}]_{ijkl}=(\bar{e}_{i}P_{R}e_{j})(\bar{e}_{k}P_{R}e_{l})$ (17)
$[O_{ee}^{S,RR}]_{1111}.$ (18)
@@
charm operators $\subset$ 4+1 flavor WET
@@
bottom, mu, tau operators $\subset$ 5+3 flavor WET
@@ 2.2 Four-fermion Operators with 5+3 Flavors
In this section, we give a list of leading heavy flavor (5+3) four-fermion operators that contribute up to the two-loop level via matching and running effects. The only new Lorentz structure appearing due to loop effects is the pure leptonic operator
@@
These two operators contribute to electron EDM at the 2-loop QED level. Otherwise, only new flavor structures are added to the EDM basis for the 5+3 theory. The complete list of relevant WET operators containing 3+1, 4+1, and 5+3 quark and lepton flavors ($n_{q}+n_{\ell}$) is given in Table 1. In this table, the operators are categorized based on their contributions to the EDMs at the LL, NLL, and NNLL order in RG improved perturbation theory. The order is indicated by the asterisk signs.
@@ 3 Electric dipole moments
Electric dipole moments are an extremely sensitive probe of $\Delta F=0$ CP-violation at low energy. The strongest constraints currently come from measurements in ThO ACME:2018yjb and HfF Roussy:2022cmp , and YbF Kara:2012ay , which can be interpreted as bounds on the electron EDM, and from bounds on the neutron EDM Abel:2020pzs and on the EDM of 199Hg Griffith:2009zz . Experiments looking for the EDM of 129Xe Sachdeva:2019rkt and 225Ra Parker:2015yka are at the moment weaker, but they provide constraints on independent combinations of hadronic parameter and Wilson coefficients Chupp:2014gka . A summary of current EDM bounds is shown in Table 2. In the future, the limits on the electron EDM are expected to improve by at least a factor of ten, while new molecular EDM experiments sensitive to nuclear Schiff moments could lead to a large improvement in the constraints on hadronic CP-violating interactions Arrowsmith-Kron:2023hcr . Several experimental collaboration, including n2EDM at the Paul Scherrer Institute n2EDM:2021yah , PanEDM Wurm:2019yfj at the Institute Laue Langevin, TUCAN EDM at Triumf Martin:2020lbx , and LANL EDM at Los Alamos National Laboratory Ito:2017ywc , aims at improving the bound on the neutron EDM by one order of magnitude. Finally, EDM of light ions such as the proton, deuteron, or 3He could be measured in upcoming storage ring experiments Alexander:2022rmq . To estimate future EDM sensitivities, we will use the projection on the second row of Table 2. In the following subsections, we will express the EDMs of charged leptons, the neutron and proton EDMs, atomic EDMs, and the CPV precession frequencies measured in molecular experiments in terms of WET operators at low energy, while in Section 4 we will show how these operators are generated from WET operators at the EW scale, via running and matching effects.
@@
expected limit $5.0\cdot 10^{-30}$ ${1.0\cdot 10^{-28}}$ $6.2\cdot 10^{-30}$ $1.0\cdot 10^{-29}$ $1.4\cdot 10^{-28}$ $1.0\cdot 10^{-27}$

### v2

# Version of June 11, 2024

@@ 1 Introduction
Because of the strength of existing and upcoming constraints, it is important to address the question of what existing EDM experiments teach us about CP violation in generic BSM scenarios. In addition, assuming multiple observations of non-zero EDMs in the next generation of experiments, we will be faced with the “inverse problem” of identifying the fundamental mechanism of CP violation from a set of low-energy measurements. Effective Field Theories (EFT) provide the theoretical tools to answer these questions with minimal reliance on specific models of BSM physics Buras:2020xsm . First of all, EFTs such as the Weak EFT (WET) Jenkins:2017jig ; Jenkins:2017dyc allow one to organize all possible flavor-diagonal ($\Delta F=0$) CP-violating operators of a given dimension. Secondly, in WET it is possible to systematically consider loop effects, such as renormalization group (RG) evolution and matching corrections induced by heavy quark thresholds, and to connect low-scale observables with physics at the electroweak scale. Here one can match WET onto other EFTs, such as the Standard Model EFT (SMEFT) Buchmuller:1985jz ; Grzadkowski:2010es or the Higgs EFT (HEFT) Buchalla:2013rka ; Buchalla:2017jlu , and study collider phenomenology. This tower of EFTs provides a general and model-independent link between EDMs, other precision flavor observables, and collider experiments, which will be necessary to solve the aforementioned inverse problem.
@@
The advantages of parameterizing EDMs in terms of a minimal set of low-energy operators have long been recognized Pospelov:2005pr ; Khriplovich:1997ga , and there is a vast literature on EDM calculations in specific BSM models (see Pospelov:2005pr ; Engel:2013lsa ; Chupp:2017rkp and references therein). In EFTs, EDMs in the SMEFT at tree level and the matching onto hadronic EFTs were considered in Ref. deVries:2012ab . Beyond the tree level, the contributions to EDMs induced by classes of SMEFT operators, including in the Yukawa Brod:2013cka ; Chien:2015xha ; Buras:2020xsm ; Brod:2022bww ; Bahl:2022yrs ; Fuchs:2020uoc ; Brod:2023wsh , top Cirigliano:2016nyn ; Fuyuto:2017xup , heavy flavor Ema:2022pmo ; Haisch:2021hcg , Higgs-gauge Dekens:2013zca ; Cirigliano:2019vfc ; Haisch:2019xyi , right-handed charged current Alioli:2017ces and flavor Endo:2019mxw ; Fajfer:2023gie sectors, have been studied in a mostly piecemeal approach. A more systematic analysis of the 1-loop contributions to EDMs in SMEFT was carried out in Ref. Kley:2021yhn , but mostly with a focus on the electron and nucleon EDMs. This paper is a first step towards a complete study of $\Delta F=0$ CP violation in EFT extensions of the SM, and its correlations with other sectors of the theory (such as the Higgs and flavor sectors). We focus here on CP-violating (CPV) $\Delta F=0$ WET operators, with $n_{f}=5$ quark flavors and $n_{\ell}=3$ lepton flavors. The main new results of this work are:
@@
  • We derive master formulae for the neutron, proton, and diamagnetic EDMs, and for the precession frequencies in molecular EDM experiments, in terms of the full set of $\Delta F=0$ WET operators at the electroweak scale. The master formulae for $\omega_{\rm HfF}$, and $\omega_{\rm ThO}$, $\omega_{\rm YbF}$ are given in Tables 7 and 15, respectively. Contributions to $d_{\rm Hg}$ are listed in Tables 8 and 9, while the neutron and proton EDMs, $d_{n}$ and $d_{p}$, are given in Tables 10 and 16, respectively.
@@
  • We consider the contributions of operators with heavy quark ($c$ and $b$) and lepton ($\mu$ and $\tau$) flavors, where renormalization group mixing and matching at the heavy fermion thresholds are crucial. In the case of scalar four-quark operators, all operators with two heavy and two light quarks contribute to EDMs at leading logarithmic accuracy. Scalar operators with four heavy quarks do not contribute at the leading log. In this case, we identify next-to-leading logarithmic contributions, arising from 1-loop running and 1-loop matching at the heavy flavor threshold. Since the renormalization group mixing is mostly driven by QCD, $d_{n}$ and $d_{\rm Hg}$ provide strong constraints even on scalar operators with only heavy quarks.
  • Four-quark vector operators with two heavy and two light quarks also contribute at the next-to-leading log. In this case, the dominant contributions arise from 2-loop running and 1-loop matching. Four-quark vector operators with two $b$ and two $c$ quarks contribute at the next-to-next-to-leading log. EDM contributions from scalar and vector operators with the same flavors are comparable.
@@
The paper is organized as follows. In Section 2 we introduce the WET basis of $\Delta F=0$ CPV operators. In Section 3 we provide expressions for the neutron, atomic, and molecular EDMs in terms of WET operators at low energy. In Section 4 we discuss the renormalization group evolution and threshold effects and in Section 5 we provide master formulae for EDMs in terms of WET coefficients at the electroweak scale. In Section 6 we discuss one example of application of the master formulae, EDM constraints on non-standard Higgs couplings. We conclude in Section 7.
@@ 2 Flavor-conserving CP-violation in WET
There are numerous sources of CP-violation in WET at the dim-4, 5, and 6 levels. Since our main goal is to study EDMs, we will restrict ourselves to flavor-conserving WET operators. We can identify two kinds of contributions. Operators constructed out of light fermion flavors, like $u,d$, or $s$ quarks or electrons, generate EDMs at the tree level. As we will see, they will receive the strongest constraints. Operators that contain heavier quark and lepton generations can also contribute via loop effects originating either from the renormalization group running or from threshold effects. Below we consider the full list of WET operators with five quark and three lepton flavors that can contribute to EDMs. Very schematically, WET Wilson coefficients at low energy are obtained by solving renormalization group equations (RGEs), which give rise to expressions of the form
@@
where $g$ here stands for any coupling in WET, and $\mu_{ew}$ is a scale close to the electroweak. The first and second terms in the bracket on the r.h.s of Eq. (1) denote the leading logarithmic (LL) and next-to-leading logarithmic (NLL) series. As we will discuss, the great majority of $\Delta F=0$ CPV operators feeds into light flavor operators at LL. These contributions are captured by the solution of the WET 1-loop RGEs, which are known Buchalla:1995vs ; Aebischer:2015fzz ; Jenkins:2017dyc . Certain classes of operators, including, for example, vector left-right operators, do not contribute at LL. In these cases, we identify NLL corrections, arising from 2-loop running and 1-loop matching at the heavy quark threshold. A summary of the order in which each WET operator contributes is given in Table 1, where operators contributing at LL, NLL, and next-to-next-to-leading log (NNLL) are indicated by no asterisk, or single and double asterisks, respectively. We stress that, while the logarithmic series expansion in Eq. (1) provides useful guidance to organize WET contributions to EDMs, and we will therefore use it throughout the paper, the RGEs involve small couplings (e.g. the QED coupling), for which the suppression from additional loops is not fully offset by the large logarithms. In these cases, it is possible that higher loop corrections to matching and running will be more important than those included here. These can only be captured by pushing the calculations of WET anomalous dimensions and matching corrections to higher order. For the sake of generality, we also include the operators that cannot be generated in SMEFT at leading order, i.e. the ones that violate the hypercharge quantum number.
In the JMS basis complete 1-loop anomalous dimension matrices (ADMs) are known Jenkins:2017jig ; Jenkins:2017dyc . Moreover, the complete 1-loop matching conditions of WET onto SMEFT has been recently worked out Dekens:2019ept in the JMS basis. Given these developments, it is useful to express the EDMs in terms of the Wilson coefficients (WCs) of the JMS basis so that their connection to UV physics can be easily established. The translation of the JMS WCs into several other popular WET bases or vice versa can be numerically carried out using the wilson Aebischer:2018bkb program, which we also use to perform the LL component of the RG running. Moreover, throughout this paper, we will follow the WCxf conventions for WET and SMEFT Aebischer:2017ugx .
@@ 2.1 Operator Basis for 3+1 Flavors
where $g_{s}$ is the strong coupling constant, $G^{A}_{\mu\nu}$ is the gluon field strength and $\widetilde{G}^{A}_{\mu\nu}=\frac{1}{2}\varepsilon_{\mu\nu\alpha\beta}G^{A\,\alpha\beta}$. Throughout this paper, we assume that the contribution of the $\bar{\theta}$ term to EDMs is canceled by the Peccei-Quinn (PQ) mechanism Peccei:1977ur ; Peccei:1977hh .
@@ 2.1.1 Dipole Operators
Here $ii$ indicates the flavor index and $T^{A}$ are the $SU(3)_{c}$ generators. Down-type type operators having $ii=11,22$, and up-type quarks with $ii=11$ directly contribute to the neutron and atomic EDMs. Dipole operators with $b$ and $c$ quarks contribute via loops. In the leptonic sector, $ii=11$, $22$, and $33$ contribute directly to electron, muon, and $\tau$ EDMs, respectively. Since the limits on $d_{\mu}$ and $d_{\tau}$ are much weaker than on the electron EDM, it is important also to consider the matching and running of heavier leptons into leptonic and semileptonic operators that contribute to molecular EDMs.
@@ 2.1.2 Weinberg Operator
The CPV Weinberg operator has the form
@@
Here $f^{ABC}$ are the structure constants of the $SU(3)_{c}$ group. Note that the WC $\big{[}{C}_{\widetilde{G}}\big{]}$ is real by definition.
@@ 2.1.3 Four-fermion operators with 3+1 Flavors
In this section, we give a complete list of four-fermion WET operators with $n_{f}=3+1$ quark and lepton flavors. All of these operators enter directly into the EDM expressions, hence these are the leading operators. A four-fermion basis in terms of scalar-pseudoscalar and tensor operators was presented in Ref. Buhler:2023gsg , where the 1-loop matching to the gradient flow scheme used in Lattice QCD calculations is also provided. Here, to facilitate the renormalization group evolution and the matching to EFTs at the electroweak scale, we adopt the JMS basis. In the first category, we consider vector operators having $(LL)(RR)$ chiralities of the fermion currents. These are given by
@@
In total, there are 24 four-quark operators with $n_{f}=3$, in agreement with Ref. Buhler:2023gsg . Semileptonic operators contribute to molecular and atomic EDMs at the tree level
@@
$c$ operators $\subset$ 4+1 flavor WET
@@
$b,\mu,\tau$ operators $\subset$ 5+3 flavor WET
@@
The relevant flavor structures for semi-leptonic operators are
$[O_{ed}^{T,RR}]_{1122}.$
$[{O}_{ee}^{S,RR}]_{ijkl}=(\bar{e}_{i}P_{R}e_{j})(\bar{e}_{k}P_{R}e_{l})$ (17)
$[O_{ee}^{S,RR}]_{1111}.$ (18)
@@ 2.2 Four-fermion Operators with 5+3 Flavors
In this section, we list the leading heavy flavor (5+3) four-fermion operators that contribute up to the 2-loop level via matching and running effects. The only new Lorentz structure appearing due to loop effects is the pure leptonic operator
@@
These two operators contribute to the electron EDM at the 2-loop QED level. Otherwise, only new flavor structures are added to the EDM basis for the 5+3 theory. The complete list of relevant WET operators containing 3+1, 4+1, and 5+3 quark and lepton flavors ($n_{q}+n_{\ell}$) is given in Table 1. In this table, the operators are categorized based on their contributions to the EDMs at the LL, NLL, and NNLL order in RG improved perturbation theory. The order is indicated by the asterisk signs.
@@ 3 Electric dipole moments
Electric dipole moments are an extremely sensitive probe of $\Delta F=0$ CP-violation at low energy. The strongest constraints currently come from measurements in ThO ACME:2018yjb , HfF Roussy:2022cmp , and YbF Kara:2012ay , which can be interpreted as bounds on the electron EDM, and from bounds on the neutron EDM Abel:2020pzs and on the EDM of 199Hg Griffith:2009zz . Experiments looking for the EDM of 129Xe Sachdeva:2019rkt and 225Ra Parker:2015yka are at the moment weaker, but they provide constraints on independent combinations of hadronic parameter and Wilson coefficients Chupp:2014gka . A summary of current EDM bounds is shown in Table 2. In the future, the limits on the electron EDM are expected to improve by at least a factor of ten, while new molecular EDM experiments sensitive to nuclear Schiff moments could lead to a large improvement in the constraints on hadronic CPV interactions Arrowsmith-Kron:2023hcr . Several experimental collaborations, including n2EDM at the Paul Scherrer Institute n2EDM:2021yah , PanEDM Wurm:2019yfj at the Institute Laue Langevin, TUCAN EDM at Triumf Martin:2020lbx , and LANL EDM at Los Alamos National Laboratory Ito:2017ywc , aim at improving the bound on the neutron EDM by one order of magnitude. Finally, EDMs of light ions such as the proton, deuteron, or 3He could be measured in storage ring experiments Alexander:2022rmq . To estimate future EDM sensitivities, we will use the projection on the second row of Table 2. In the following subsections, we will express the EDMs of charged leptons, the neutron and proton EDMs, atomic EDMs, and the CPV precession frequencies measured in molecular experiments in terms of WET operators at low energy, while in Section 4 we will show how these operators are generated from WET operators at the EW scale, via running and matching effects.
@@
expected limit $4.0\cdot 10^{-29}$ ${1.0\cdot 10^{-28}}$ $6.2\cdot 10^{-30}$ $1.0\cdot 10^{-29}$ $1.4\cdot 10^{-28}$ $1.0\cdot 10^{-27}$